# Information Physics of Intelligence: Unifying Logical Depth and Entropy under Thermodynamic Constraints

### v2

@@ IV-A Landauer’s Principle and Energy Bounds
Deriving formula $\varphi$ from knowledge base $K$ requires minimum energy:
$E_{\text{compute}}(\varphi|K)\geq\text{Ld}(\varphi|K)\cdot k_{B}T\ln(2)$ (71)
@@
Proof.
Each inference step in deriving $\varphi$ from $K$ performs an irreversible logical operation. By Landauer’s principle [landauer1961irreversibility], each irreversible operation erasing or transforming 1 bit of information requires minimum energy $k_{B}T\ln(2)$.
Consider a derivation with $\text{Ld}(\varphi|K)$ steps. Each step:
@@
The total bit erasures across all steps is at least $\text{Ld}(\varphi|K)$. By Landauer’s principle, total energy:
$E_{\text{compute}}\geq\text{Ld}(\varphi|K)\cdot k_{B}T\ln(2)$ (72)
@@ IV-B Energy-Time-Space Triality Bound
For any hybrid strategy $\Pi=(S,C,R)$ answering query distribution $P_{Q}$ over knowledge base $K$ in the information-rich regime:
$E(\Pi)\cdot T(\Pi)\cdot M^{-1}(\Pi)\geq\Omega(H(Q|K))\cdot k_{B}T\ln(2)$ (74)
where $\Omega(H(Q|K))$ represents the computational cost induced by the conditional entropy $H(Q|K)$, with dimension of time. Via Landauer’s principle, both sides can be dimensionally reduced to time, ensuring consistency.
@@
$H(Q|K)\gg\log_{2}(|\text{Atom}(K)|),\log_{2}(Ld(Q|K))$ (75)
By Corollary 1, this implies:
$Ld(Q|K)\cdot\ln(2)\approx H(Q|K)\quad[\text{asymptotically}]$ (76)
@@
$E(\Pi_{c})=\mathbb{E}_{q\sim P_{Q}}[E_{\text{compute}}(q|K)]\geq Ld(Q|K)\cdot k_{B}T\ln(2)$ (78)
@@
$\begin{split}E(\Pi_{c})\cdot T(\Pi_{c})&\geq(Ld(Q|K))^{2}\cdot k_{B}T\ln(2)\\ &\geq\Omega(H(Q|K)^{2}/\ln(2))\cdot k_{B}T\ln(2)\\ &=\Omega(H(Q|K)^{2}\cdot k_{B}T)\end{split}$ (80)
@@
Setup: - Precompute all answers: $S=\{(q,A(q,K)):q\in\text{supp}(P_{Q})\}$. - Storage requirement:
$M(\Pi_{s})=H(Q|K)$ (81)
@@
$E(\Pi_{s})=H(Q|K)\cdot k_{B}T\ln(2)\cdot t_{\text{avg}}\cdot t_{\text{refresh}}$ (82)
@@
Analysis: By the Storage-Computation Duality Theorem (Theorem 4):
$\min[\text{Storage}]+\min[\text{Computation}]\geq H(Q|K)\cdot(1+\ln(2))$ (85)
@@
$E(\Pi_{h})\cdot T(\Pi_{h})\cdot M^{-1}(\Pi_{h})\geq\Omega(H(Q|K)\cdot k_{B}T\ln(2))$ (86)
@@
$M_{\text{critical}}=\frac{H(Q)}{\log\left(\frac{E_{\text{budget}}}{H(Q|K)\cdot k_{B}T\ln(2)}\right)}$ (87)
@@ V-A Threshold Policies for Hybrid Strategies
Theorem 4 implies that the marginal benefit of storing additional information diminishes as the knowledge base grows. For a query $q$ with access frequency $f_{q}$, the expected cost of computing on-demand is
$\mathbb{E}[\text{Compute Cost}]=f_{q}\cdot H_{\mathrm{derive}}(q\mid K)\\ =\Theta\!\left(f_{q}\cdot\frac{H_{q}}{\log|\text{Atom}(K)|}\right),$ (91)
@@
$f_{q}\cdot\frac{H_{q}}{\log|\text{Atom}(K)|}>H_{q}\quad\Longleftrightarrow\quad f_{q}>\Theta(\log|\text{Atom}(K)|).$ (93)
@@
$f_{q}\cdot\frac{L_{d}(q\mid K)}{H_{q}}>\tau,$ (94)
where the threshold $\tau=\Theta(\log|\text{Atom}(K)|)$ grows logarithmically with knowledge base size.
This policy generalizes classical caching (e.g., LFU) by incorporating computational depth $L_{d}(q|K)$ and semantic information $H_{q}$, and explains the phase transition in storage utility observed empirically.
@@ V-B Phase Transition
$\alpha_{c}=\frac{H(Q)}{\mathbb{E}[\text{Ld}(Q|K)]\cdot\ln(2)}$ (95)
@@ VI Numerical Validation and Simulations
All datasets satisfy $H(q\mid K)\gg\log|\mathrm{Atom}(K)|$ (verified empirically), placing them firmly in the information-rich regime. We validate our theoretical predictions through extensive numerical experiments on synthetic knowledge systems. Knowledge bases range from 1K to 100K atoms. Query distributions include uniform and Zipfian ($\alpha\in\{1.0,1.2,1.5,1.8,2.0\}$). Physical parameters are: temperature T=300T=300K.
In Experiment 1, we validate Theorem 4 (Frequency-Weighted Duality) by testing individual queries sampled from realistic distributions. For each query $q$, we compute its access frequency $f_{q}$ from a Zipfian distribution, then verify that the optimal amortized cost satisfies the theoretical bound $\text{Cost}_{\text{amort}}(q)\geq H_{q}(1+\frac{1}{c\log|K|}-\frac{1}{f_{q}})$. We test 50 randomly sampled queries from each of three knowledge bases: Small (1K atoms), Medium (10K atoms), Large (100K atoms), with average logical depths 3–5.
TABLE I: Single-Query Duality Theorem Validation (Theorem 4)
KB Size Entities Queries Tested Satisfaction Rate Avg Margin
Small 1.00K 50 100.0% 6785.725
Medium 10.00K 50 100.0% 280828.276
Large 100.00K 50 100.0% 10329841.734
Note: Satisfaction Rate = fraction of queries satisfying $\text{Cost}_{\text{amort}}(q)\geq H_{q}(1+\frac{1}{c\log|K|}-\frac{1}{f_{q}})$. Avg Margin = average difference between observed cost and theoretical bound.
@@
Physical Interpretation: At room temperature with $H(Q|K)=10^{6}$ bits, the fundamental limit is:
@@
TABLE II: Baseline Comparison at Cache Size = 50 (5% of queries)
@@ VII-A Large Language Models
Consider a GPT-4 scale model with 1.7T parameters. The knowledge base contains $H(K)=10^{15}$ bits, and queries follow a power-law distribution with $H(Q)=10^{12}$ bits.
Parameter Settings. Average logical depth: $\mathbb{E}[\text{Ld}]=5$ inference steps. Energy budget: $E_{\text{budget}}=1000$ kWh. Temperature: T=300T=300K.
@@
$M_{\text{critical}}=\frac{10^{12}}{\log(1000\times 3.6\times 10^{9}/(10^{6}\times k_{B}T\ln 2))}\approx 2.5\times 10^{11}\text{ bits}$ (97)
Optimal Strategy. Following our threshold policy framework, we compute $\text{Ld}^{*}=2.5$ and $f^{*}=0.85$. The optimal strategy:
  • Parametric storage (70%): Facts with $\text{Ld}\leq 2$ and frequency >0.85>0.85
  • RAG retrieval (25%): Facts with $3\leq\text{Ld}\leq 8$
  • Dynamic reasoning (5%): Complex queries with $\text{Ld}>8$
@@ VII-B Knowledge Graphs
Query Type Ld Frequency Example
@@ VII-C Edge Intelligence
Hybrid Strategy. Device storage (200 MB): Common sense, personal habits, quick facts. Cloud coordination: Low-frequency knowledge and complex reasoning ($\text{Ld}>10$). Adaptive caching: Weekly updates based on usage patterns.
@@ VIII-A Information Theory and Kolmogorov Complexity
Bennett’s logical depth [bennett1988logical] addresses this gap by measuring computational sophistication: $\text{Ld}(x)$ is the runtime of the fastest program generating $x$ from a short input. This distinguishes between ”shallow” information (random bits, easy to describe but not to compute from scratch) and ”deep” information (organized structures requiring significant computation to generate).
@@
    Conditional depth: We formalize $\text{Ld}(\varphi|\psi)$ for conditional derivation, enabling analysis of incremental inference.
@@
    Information-theoretic duality: We prove that derivation entropy $H_{\text{derive}}=\text{Ld}\cdot\ln(2)$ is dual to Shannon entropy, establishing a precise quantitative relationship.
@@ VIII-B Thermodynamics of Computation
    Triality bound: We derive the first joint bound on energy, time, and storage: $E\cdot T\cdot M^{-1}\geq\Omega(H(Q|K))\cdot k_{B}T\ln(2)$, where $\Omega(H(Q|K))$ represents computational cost with time dimension. Previous work bounded resources independently; our triality captures fundamental coupling.
@@ VIII-F Large Language Models and Retrieval
Our analysis reveals that optimal LLM architectures should store high-frequency, high-logical-depth knowledge parametrically, while retrieving low-frequency or simple facts. This explains the empirical success of RAG: it exploits the storage-computation duality by allocating expensive parametric storage to knowledge with high $\text{Ld}(q|K)\times\text{freq}(q)$ product.
@@ VIII-H Phase Transitions in Computer Science
The transition we observe is sharp and universal across problem domains, determined solely by $H(Q)$, $\mathbb{E}[\text{Ld}]$, and physical constants. This universality distinguishes it from problem-specific phase transitions in SAT solvers or random graphs.
@@ VIII-I Information-Theoretic Learning and Prediction
The connection arises through regret minimization: our storage selection problem is equivalent to minimizing regret in an online learning game where ”experts” are candidate storage sets. However, standard online learning assumes unit costs per expert consultation, whereas in our setting, computation cost scales with logical depth $\text{Ld}(q|S)$, requiring depth-aware algorithms.
@@ IX Conclusion
In this paper, we have established a unified physical framework for intelligence, grounded in the fundamental axiom that information constitutes the enabling mapping from ontological states to carrier states [xu2025information, xu2024research]. By rigorously analyzing the thermodynamic costs associated with this mapping, we have demonstrated that the separation between ”static storage” (Shannon Entropy) and ”dynamic computation” (Logical Depth) is artificial; they are, in fact, conjugate variables governed by a fundamental constraint: the Energy-Time-Space Triality Bound.
@@ Acknowledgment
During the writing and revision of this paper, we received many insightful comments from Associate Professor Wang Rui of the School of Computer Science at Shanghai Jiao Tong University and also gained much inspiration and assistance from regular academic discussions with doctoral students Yiming Wang, Chun Li, Hu Xu, Siyuan Qiu, Junxuan He, and Xiao Wang. We hereby express our sincere gratitude to them.

### v3

@@ IV-A Landauer’s Principle and Energy Bounds
Under the assumption that each inference step constitutes an ideal logically irreversible operation, deriving formula $\phi$ from knowledge base $S_{O}$ requires minimum energy:
$E_{\text{compute}}(\varphi|S_{O})\geq\text{Dd}(\varphi|S_{O})\cdot k_{B}T\ln(2)$ (72)
@@
Each inference step in deriving $\varphi$ from $S_{O}$ performs an irreversible logical operation. By Landauer’s principle [landauer1961irreversibility], each irreversible operation erasing or transforming 1 bit of information requires minimum energy $k_{B}T\ln(2)$.
Consider a derivation with $\text{Dd}(\varphi|S_{O})$ steps. Each step:
@@
The total bit erasures across all steps is at least $\text{Dd}(\varphi|S_{O})$. By Landauer’s principle, total energy:
$E_{\text{compute}}\geq\text{Dd}(\varphi|S_{O})\cdot k_{B}T\ln(2)$ (73)
@@ IV-B Energy-Time-Space Triality Bound
For any hybrid strategy $\Pi=(S,C,R)$ answering query distribution $P_{Q}$ over knowledge base $S_{O}$ in the information-rich regime:
$E(\Pi)\cdot T(\Pi)\cdot M^{-1}(\Pi)\geq\Omega(H(Q|S_{O})\cdot k_{B}T\ln(2))$ (75)
@@
$H(Q|S_{O})\gg\log_{2}(|\text{Atom}(S_{O})|),\log_{2}(Dd(Q|S_{O}))$ (76)
By Corollary LABEL:cor:Asymptotic_Equivalence, this implies:
$Dd(Q|S_{O})\cdot\ln(2)\approx H(Q|S_{O})\quad[\text{asymptotically}]$ (77)
@@
$E(\Pi_{c})=\mathbb{E}_{q\sim P_{Q}}[E_{\text{compute}}(q|S_{O})]\geq Dd(Q|S_{O})\cdot k_{B}T\ln(2)$ (79)
@@
$\begin{split}E(\Pi_{c})\cdot T(\Pi_{c})&\geq(Dd(Q|S_{O}))^{2}\cdot k_{B}T\ln(2)\\ &\geq\Omega(H(Q|S_{O})^{2}/\ln(2))\cdot k_{B}T\ln(2)\\ &=\Omega(H(Q|S_{O})^{2}\cdot k_{B}T)\end{split}$ (81)
@@
Setup: - Precompute all answers: $S=\{(q,A(q,S_{O})):q\in\text{supp}(P_{Q})\}$. - Storage requirement:
$M(\Pi_{s})=H(Q|S_{O})$ (82)
@@
$E(\Pi_{s})=H(Q|S_{O})\cdot k_{B}T\ln(2)\cdot t_{\text{avg}}\cdot t_{\text{refresh}}$ (83)
@@
Analysis: By the Storage-Computation Duality Theorem (Theorem LABEL:thm:duality):
$\min[\text{Storage}]+\min[\text{Computation}]\geq H(Q|S_{O})\cdot(1+\ln(2))$ (86)
@@
$E(\Pi_{h})\cdot T(\Pi_{h})\cdot M^{-1}(\Pi_{h})\geq\Omega(H(Q|S_{O})\cdot k_{B}T\ln(2))$ (87)
@@
$M_{\text{critical}}=\frac{H(Q)}{\log\left(\frac{E_{\text{budget}}}{H(Q|S_{O})\cdot k_{B}T\ln(2)}\right)}$ (88)
@@ V-A Threshold Policies for Hybrid Strategies
Theorem LABEL:thm:duality implies that the marginal benefit of storing additional information diminishes as the knowledge base grows. For a query $q$ with access frequency $f_{q}$, the expected cost of computing on-demand is
$\mathbb{E}[\text{Compute Cost}]=f_{q}\cdot H_{\mathrm{derive}}(q\mid S_{O})\\ =\Theta\!\left(f_{q}\cdot\frac{H_{q}}{\log|\text{Atom}(S_{O})|}\right),$ (92)
@@
$f_{q}\cdot\frac{H_{q}}{\log|\text{Atom}(S_{O})|}>H_{q}\quad\Longleftrightarrow\quad f_{q}>\Theta(\log|\text{Atom}(S_{O})|).$ (94)
@@
$f_{q}\cdot\frac{Dd(q\mid S_{O})}{H_{q}}>\tau,$ (95)
where the threshold $\tau=\Theta(\log|\text{Atom}(S_{O})|)$ grows logarithmically with knowledge base size.
This policy generalizes classical caching (e.g., LFU) by incorporating computational depth $Dd(q|S_{O})$ and semantic information $H_{q}$, and explains the phase transition in storage utility observed empirically.
@@ V-B Phase Transition
$\alpha_{c}=\frac{H(Q)}{\mathbb{E}[\text{Dd}(Q|S_{O})]\cdot\ln(2)}$ (96)
@@ VI Numerical Validation and Simulations
All datasets satisfy $H(q\mid S_{O})\gg\log|\mathrm{Atom}(S_{O})|$ (verified empirically), placing them firmly in the information-rich regime. We validate our theoretical predictions through extensive numerical experiments on synthetic knowledge systems. Knowledge bases range from 1K to 100K atoms. Query distributions include uniform and Zipfian ($\alpha\in\{1.0,1.2,1.5,1.8,2.0\}$). Physical parameters are: temperature T=300T=300K.
In Experiment 1, we validate Theorem LABEL:thm:duality (Frequency-Weighted Duality) by testing individual queries sampled from realistic distributions. For each query $q$, we compute its access frequency $f_{q}$ from a Zipfian distribution, then verify that the optimal amortized cost satisfies the theoretical bound $\text{Cost}_{\text{amort}}(q)\geq H_{q}(1+\frac{1}{c\log|K|}-\frac{1}{f_{q}})$. We test 50 randomly sampled queries from each of three knowledge bases: Small (1K atoms), Medium (10K atoms), Large (100K atoms), with average logical depths 3–5.
TABLE II: Single-Query Duality Theorem Validation (Theorem LABEL:thm:frequency_weighted_duality)
KB Size Entities # Queries Sat. Rate Avg. Margin
Small 1.00K 50 100.0% 6,785.7
Medium 10.00K 50 100.0% 280,828.3
Large 100.00K 50 100.0% 10,329,841.7
Note: Sat. Rate = fraction of queries satisfying $\mathrm{Cost}_{\mathrm{amort}}(q)\geq H_{q}\bigl(1+\frac{1}{c\log|K|}-\frac{1}{f_{q}}\bigr)$. Avg. Margin = average difference between observed cost and theoretical bound.
@@
Physical Interpretation: At room temperature with $H(Q|S_{O})=10^{6}$ bits, the fundamental limit is:
@@
TABLE III: Baseline Comparison at Cache Size = 50 (5% of queries)
@@ VII-A Large Language Models
Consider a GPT-4 scale model with 1.7T parameters. The knowledge base contains $H(S_{O})=10^{15}$ bits, and queries follow a power-law distribution with $H(Q)=10^{12}$ bits.
Parameter Settings. Average logical depth: $\mathbb{E}[\text{Dd}]=5$ inference steps. Energy budget: $E_{\text{budget}}=1000$ kWh. Temperature: T=300T=300K.
@@
$M_{\text{critical}}=\frac{10^{12}}{\log(1000\times 3.6\times 10^{9}/(10^{6}\times k_{B}T\ln 2))}\\ \approx 2.5\times 10^{11}\text{ bits}$ (98)
Optimal Strategy. Following our threshold policy framework, we compute $\text{Dd}^{*}=2.5$ and $f^{*}=0.85$. The optimal strategy:
  • Parametric storage (70%): Facts with $\text{Dd}\leq 2$ and frequency >0.85>0.85
  • RAG retrieval (25%): Facts with $3\leq\text{Dd}\leq 8$
  • Dynamic reasoning (5%): Complex queries with $\text{Dd}>8$
@@ VII-B Knowledge Graphs
TABLE IV: Query Patterns in the Biomedical KB
Query Type Dd Freq. Example
@@ VII-C Edge Intelligence
Hybrid Strategy. Device storage (200 MB): Common sense, personal habits, quick facts. Cloud coordination: Low-frequency knowledge and complex reasoning ($\text{Dd}>10$). Adaptive caching: Weekly updates based on usage patterns.
@@ VIII-A Information Theory and Kolmogorov Complexity
Bennett’s logical depth [bennett1988logical] addresses this gap by measuring computational sophistication: $\text{Dd}(x)$ is the runtime of the fastest program generating $x$ from a short input. This distinguishes between ”shallow” information (random bits, easy to describe but not to compute from scratch) and ”deep” information (organized structures requiring significant computation to generate).
@@
    Conditional depth: We formalize $\text{Dd}(\varphi|\psi)$ for conditional derivation, enabling analysis of incremental inference.
@@
    Information-theoretic duality: We prove that derivation entropy $H_{\text{derive}}=\text{Dd}\cdot\ln(2)$ is dual to Shannon entropy, establishing a precise quantitative relationship.
@@ VIII-B Thermodynamics of Computation
    Triality bound: We derive the first joint bound on energy, time, and storage: $E\cdot T\cdot M^{-1}\geq\Omega(H(Q|S_{O})\cdot k_{B}T\ln(2))$. Previous work bounded resources independently; our triality captures fundamental coupling.
@@ VIII-F Large Language Models and Retrieval
Our analysis reveals that optimal LLM architectures should store high-frequency, high-logical-depth knowledge parametrically, while retrieving low-frequency or simple facts. This explains the empirical success of RAG: it exploits the storage-computation duality by allocating expensive parametric storage to knowledge with high $\text{Dd}(q|S_{O})\times\text{freq}(q)$ product.
@@ VIII-H Phase Transitions in Computer Science
The transition we observe is sharp and universal across problem domains, determined solely by $H(Q)$, $\mathbb{E}[\text{Dd}]$, and physical constants. This universality distinguishes it from problem-specific phase transitions in SAT solvers or random graphs.
@@ VIII-I Information-Theoretic Learning and Prediction
The connection arises through regret minimization: our storage selection problem is equivalent to minimizing regret in an online learning game where ”experts” are candidate storage sets. However, standard online learning assumes unit costs per expert consultation, whereas in our setting, computation cost scales with logical depth $\text{Dd}(q|S)$, requiring depth-aware algorithms.
@@ IX Conclusion
In this paper, we have established a unified physical framework for intelligence, grounded in the fundamental assumption that information constitutes the enabling mapping from ontological states to carrier states [xu2025information, xu2024research]. By rigorously analyzing the thermodynamic costs associated with this mapping, we have demonstrated that the separation between ”static storage” (Shannon Entropy) and ”dynamic computation” (Logical Depth) is artificial; they are, in fact, conjugate variables governed by a fundamental constraint: the Energy-Time-Space Triality Bound.
@@ Acknowledgment
During the writing and revision of this paper, we received many insightful comments from Associate Professor Wang Rui of the School of Computer Science at Shanghai Jiao Tong University and also gained much inspiration and assistance from regular academic discussions with doctoral students Yiming Wang, Chun Li, Hu Xu, Siyuan Qiu, Jiashuo Zhang, Junxuan He, and Xiao Wang. We hereby express our sincere gratitude to them.